# Invariants for $\mathbb{G}_{(r)}$-modules

### v1

@@ 6. Vector bundles
Let $(\mathbb{G},\mathcal{E}_{(r)}$ be an affine group scheme of exponential type of height $r$. Consider the bigraded algebra $k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]$ where both $s\otimes 1$ and $t\otimes 1$ are given the bigrading $(1,0)$. We define $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}$ to be
@@
$\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}\quad\equiv\quad(s\otimes 1)(1\otimes\Theta_{\mathfrak{g},r}^{\exp})\ +(1\otimes t)(1\otimes\Phi_{\mathfrak{g},r}^{*}(\Theta_{\mathbb{G}_{(r)}}))\ \in\ k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]\otimes k\mathbb{G}_{(r)}.$
@@
So defined, $\Theta_{\mathbb{G}/\mathbb{P}^{1},r}$ is bihomogenous of bidegree $(1,p^{r-1})$, thereby determining a global section of $\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}(1,p^{r-1})$.
Remark 6.2.
Specializing Definition 6.1 to $\mathbb{G}=\mathbb{G}_{a}$, we see by Example 2.10 that $\Theta_{\mathbb{G},r}^{exp}$ can be identified with $\Theta_{\mathbb{G}_{a}^{\times r},1}$ as elements of $k[x_{0},\ldots,x_{r-1}]\otimes k[v_{0},\ldots,v_{r-1}]/(v_{i}^{p})$.
@@
$JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \frac{m}{p}[p]\ \ \iff\ \ JT(\Theta_{\mathfrak{g}/\mathbb{P}^{1},r,(\underline{B},\langle s,t\rangle)},k(\underline{B},\langle a,b\rangle\otimes M)\ =\ \frac{m}{p}[p].$
@@
As recalled in the proof of Proposition 4.10, $\Theta_{\mathfrak{g},r,\underline{B}}-\Theta_{\mathfrak{g},r,\underline{B}}^{\exp}$ is a linear combination of terms each of which is a product of at least 2 $p$-nilpotent factors. Thus, proposition follows from [6, Prop 8]. ∎
In the following proposition, we utilize $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathfrak{g})\to V_{r}(\mathbb{G})$, the bicontinuous morphism of Lemma 4.7, and the construction in [14, Prop 4.8] of the quasi-coherent sheaf ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathbb{G},r,M})^{j}\})$ on $\mathbb{P}V_{r}(\mathfrak{g})$. As seen [14], these kernel sheaves (as well as cokernel and image sheaves) are algebraic vector bundles provided that $M$ has constant Jordan type.
@@
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$, and let $M$ be a $\mathbb{G}_{(r)}$-module. Then for any $j,0<j<p$, there is a quasi-coherent sheaf ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ on $\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$ satisfying
@@
    The fiber over $\langle 0,t\rangle\in\mathbb{P}^{1}$ of ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ is the quasi-coherent sheaf $\Phi_{\mathfrak{g},r}^{*}({\mathcal{K}er}\{(\tilde{\Theta}_{\mathbb{G},r,M})^{j}\})$, the pull-back along $\Phi_{\mathfrak{g},r}$ of the quasi-coherent sheaf associated to the graded $k[V_{r}(\mathfrak{g})]$-module ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{j}:k[V_{r}(\mathfrak{g})]\otimes M\to(k[V_{r}(\mathfrak{g})]\otimes M)(jp^{r-1})\}.$
@@
    The fiber over $\langle s,0\rangle\in\mathbb{P}^{1}$ of ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ is the quasi-coherent sheaf associated to the graded $k[\mathcal{C}_{r}(\mathfrak{g})]$-module
    $ker\{(\Theta_{\mathfrak{g},r,M}^{exp})^{j}:k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M\to(k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M)(jp^{r-1})\}.$
@@
    If $M$ is a finite dimensional $\mathbb{G}_{(r)}$-module of constant Jordan type, then ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ is a locally free, coherent sheaf which restricts to the locally free coherent sheaves $\Phi_{\mathfrak{g},r}^{*}({\mathcal{K}er}\{(\tilde{\Theta}_{\mathbb{G},r,M})^{j}\}),\ {\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\}$ on $\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$.
In other words, for any $\mathbb{G}_{(r)}$-module of constant Jordan type, taking kernels of $(\tilde{\Theta}_{\mathbb{G},r,M})^{j}$ and $(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}$ yield $\mathbb{A}^{1}$-equivalent vector bundles on $\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$.
A completely parallel construction applies to taking images of images and cokernels of $(\tilde{\Theta}_{\mathbb{G},r,M})^{j}$ and $(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}$.
@@
We consider the bigraded algebra $k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]$ where both $s\otimes 1$ and $t\otimes 1$ are given the bigrading $(1,0)$. We define $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}$ to be
$\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}\quad\equiv\quad(s\otimes 1)(1\otimes\Theta_{\mathfrak{g},r}^{\exp})\ +(1\otimes t)(1\otimes\Phi_{\mathfrak{g},r}^{*}(\Theta_{\mathbb{G}_{(r)}}))\ \in\ k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]\otimes k\mathbb{G}_{(r)}.$
@@
By Theorem 4.8(1), $\Theta_{\mathfrak{g},r}=\Phi_{\mathfrak{g},r}^{*}(\Theta_{\mathbb{G}_{(r)}})$. By Proposition 4.5 and Theorem 4.8, $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}$ is bihomogeneous of bidegree $(1,p^{r-1})$. Thus, the bigraded module
(6.4.1) $ker\{(\Theta_{\mathfrak{g}/\mathbb{P}^{1},r})^{j}:(k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M)\ \to\ (k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M)[1,p^{r-1}]$
@@
determines the quasi-coherent sheaf ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ on $\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$. The first two assertions follow from the observation that the fiber above $\langle a,b\rangle\in\mathbb{P}^{1}$ of ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ is the sheaf associated to the bigraded graded module (6.4.1) obtained by extending along $k[s,t]\to k$ sending $s$ to $a$ and $t$ to $b$.

### v2

@@ 6. Vector bundles and stratifications
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$, let $r$ be a chosen positive integer, $j$ and integer with $1\leq j<r$, and let $M$ be a $\mathbb{G}_{(r)}$-module. We define the following coherent sheaves on $\mathbb{P}V_{r}(\mathbb{G})$
$\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\quad\mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\quad\mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\}$
as the sheaves associated to kernel, cokernel, and image of the $j$-th iterate of the map of graded $k[V_{r}(\mathbb{G})]$-modules
$\tilde{\Theta}_{\mathbb{G},r,M}:(k[V_{r}(\mathbb{G})]\otimes M)^{\sim}\quad\to\quad(k[V_{r}(\mathbb{G})]\otimes M)^{\sim}(p^{r-1})$
given in Theorem 2.18 (1).
Definition 6.2.
@@
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$, $j$ and integer with $1\leq j<r$, and let $M$ be a $\mathbb{G}_{(r)}$-module. We define the following coherent sheaves on $\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$
$\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathfrak{g},r,M}^{j}\},\quad\mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathfrak{g},r,M}^{j}\},\quad\mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathfrak{g},r,M}^{j}\}$
@@
$\mathcal{K}\text{\it er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\quad\mathcal{C}\text{\it oker}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}\},\quad\mathcal{I}\text{\it m}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}\}$
as the sheaves associated to kernel, cokernel, and image of the $j$-th iterate of the maps $\tilde{\Theta}_{\mathfrak{g},r,M}$ and $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}$ of graded $k[\mathcal{C}_{r}(\mathfrak{g})]$-modules given in Proposition 4.11 (1) and Proposition 4.17 (1).
We next verify that the coherent sheaves of Definition 6.2 are locally constant on each stratum.
Theorem 6.3.
Let $\mathbb{G}$ be an affine group scheme of finite type of $k$ and let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module. Then for each Jordan type $\underline{a}\ =\ \sum_{i=1}^{p}a_{i}[i]$, the restrictions of the coherent sheaves $\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\}$ to the locally closed subvariety $JT_{\mathbb{G},r,M}^{-1}(\underline{a})\ \subset\ \mathbb{P}V_{r}(\mathbb{G})$ are locally free coherent sheaves on $JT_{\mathbb{G},r,M}^{-1}(\underline{a})$.
If $(\mathbb{G},\mathcal{E}_{(r)})$ is an affine group scheme of exponential type of height $r$, then the pull-backs along $\Phi_{\mathfrak{g},r}:JT_{\mathfrak{g},r,M}^{-1}(\underline{a})\to JT_{\mathbb{G},r,M}^{-1}(\underline{a})$ of the sheaves $ker\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\\ coker\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ im\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\}$ are locally free sheaves on $JT_{\mathfrak{g},r,M}^{-1}(\underline{a})\ \subset\ \mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$.
Moreover, if $(\mathbb{G},\mathcal{E}_{(r)})$ is an affine group scheme of exponential type of height $r$, then the restrictions of the coherent sheaves $\mathcal{K}\text{\it er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\ \mathcal{C}\text{\it oker}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\\ \mathcal{I}\text{\it m}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\}$ to $(JT_{\mathfrak{g},r,M}^{exp})^{-1}(\underline{a})\ \subset\ \mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$ are locally free sheaves for any Jordan type $\underline{a}\ =\ \sum_{i=1}^{p}a_{i}[i]$.
Proof.
Let $Y_{\underline{a}}$ denote $(JT_{\mathfrak{g},r,M}^{exp})^{-1}(\underline{a})\ \subset\ \mathbb{P}V_{r}(\mathfrak{g})$ and consider its closure
$\overline{Y_{\underline{a}}}\ \equiv\ JT_{\mathfrak{g},r,M}^{-1}(\overline{\{\underline{a}\}})\ =\ \bigcup_{\underline{b}\leq\underline{a}}JT_{\mathfrak{g},r,M}^{-1}(\underline{b})\quad\hookrightarrow\quad\mathbb{P}V_{r}(\mathbb{G})$
corresponding to the homogeneous ideal $I_{\underline{a}}\ \subset\ k[V_{r}(\mathbb{G})]$. Consider $k[\overline{Y_{\underline{a}}}]\ =\ k[V_{r}(\mathbb{G})]/I_{\underline{a}}$, equipped with the grading inherited from the grading of $k[V_{r}(\mathbb{G})]$ and set
$\Theta_{\mathbb{G},r;\overline{Y_{\underline{a}}}}\ \equiv\ k[\overline{Y_{\underline{a}}}]\otimes_{k[V_{r}(\mathbb{G})]}\Theta_{\mathbb{G},r}.$
So defined, $\Theta_{\mathbb{G},r;\overline{Y_{\underline{a}}}}$ determines a $p$-nilpotent $k[\overline{Y_{\underline{a}}}]$-linear operator $\Theta_{\mathbb{G},r,M:\overline{Y_{\underline{a}}}}$ on $k[\overline{Y_{\underline{a}}}]\otimes M$ as in Theorem 2.15. As in Theorem 2.18, $\Theta_{\mathbb{G},r;\overline{Y_{\underline{a}}}}$ determines
$\tilde{\Theta}_{\mathbb{G},r;\overline{Y_{\underline{a}}}}\ \in\ \Gamma(\overline{Y_{\underline{a}}},\mathcal{O}_{\overline{Y_{\underline{a}}}}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
and the map of graded $k[\overline{Y_{\underline{a}}}]$-modules
$\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}:k[\overline{Y_{\underline{a}}}]\otimes M)^{\sim}\ \to\ (k[\overline{Y_{\underline{a}}}]\otimes M)^{\sim}(p^{r-1}).$
Then the restrictions of $\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\}$ along $Y_{\underline{a}}\subset\overline{Y_{\underline{a}}}\ \hookrightarrow\ \mathbb{P}V_{r}(\mathbb{G})$ are identified with the restrictions to the open subset $Y_{\underline{a}}\subset\overline{Y_{\underline{a}}}$ of $\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\},\ \mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\},\ \mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\}$.
The proof that the restrictions to $Y_{\underline{a}}$ of $\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\},\ \mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\}$ and $\mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\}$ are locally free coherent sheaves on $Y_{\underline{a}}$ is completed by repeating the proof of [16, Thm 4.13] with $U\subset\mathbb{P}(G),\ \tilde{\Theta}_{G}$ of that theorem replaced by $Y_{\underline{a}}\ \subset\ \overline{Y_{\underline{a}}},\ \tilde{\Theta}_{\mathbb{G},r;\overline{Y_{\underline{a}}}}$ as above.
If $(\mathbb{G},\mathcal{E}_{(r)})$ is an affine group scheme of exponential type of height $r$, then the pull-backs of these locally free sheaves along along the homeomoprhism $\Phi_{\mathfrak{g},r}:JT_{\mathfrak{g},r,M}^{-1}(\underline{a})\to JT_{\mathbb{G},r,M}^{-1}(\underline{a})$ are also locally free. Finally, only notational changes replacing $\Theta_{\mathfrak{g},r,M}$ by $\Theta_{\mathfrak{g},r,M}^{exp}$ are needed to obtains proofs of the assertions for $\mathcal{K}\text{\it er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\ \mathcal{C}\text{\it oker}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\mathcal{I}\text{\it m}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\}$. ∎
Example 3.1 gives examples of $\mathbb{G}_{(r)}$ modules leading to a stratification of $\mathbb{P}V_{r}(\mathbb{G})$ with a single stratum and Example 3.2 gives examples of $\mathbb{G}_{(r)}$-modules with exactly two strata.
We give one “trivial” example of a $\mathbb{G}_{(r)}$ module with more than two strata, primarily to show that some examples can be easily computed.
Example 6.4.
Let $\mathbb{G}=\mathbb{G}_{a}^{\times p}$, so that $k\mathbb{G}_{(1)}=k[u_{0},\dots,u_{p-1}]$. Consider the $k\mathbb{G}_{(1)}$-module $M$ defined as the quotient of $k\mathbb{G}_{(1)}$ by the ideal generated by $\{u_{i}^{i}\}$. Consider points $\underline{b}=\langle b_{0},\ldots,b_{p-1}\rangle\in\mathbb{P}V_{1}(\mathbb{G})\simeq\mathbb{P}^{p-1}$. If $\underline{b}$ is given by $b_{i}=1,\ b_{j}=0,j\not=i$, then $JT_{\mathbb{G},1,M}(\underline{b})$ equals $\sum_{j\not=i}[j]+i[1]$.
We next investigate the relationship between the coherent sheaves $\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathfrak{g},r,M}^{j}\},\\ \mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\}$ and the coherent sheaves $\mathcal{K}\text{\it er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\\ \mathcal{C}\text{\it oker}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}\},\ \mathcal{I}\text{\it m}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}\}$.
Definition 6.5.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. Consider the bigraded algebra $k[s,t]\otimes k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$ where both $s\otimes 1$ and $t\otimes 1$ are given the bigrading $(1,0)$. We define $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}$ to be
(6.5.1) $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}\quad\equiv\quad(s\otimes 1)(1\otimes\Theta_{\mathfrak{g},r}^{\exp})\ +(t\otimes 1)(1\otimes\Phi_{\mathfrak{g},r}^{*}(\Theta_{\mathbb{G}_{(r)}}))\ \in\ k[s,t]\otimes k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes k\mathbb{G}_{(r)}.$
@@
So defined, $\Theta_{\mathbb{G}/\mathbb{P}^{1},r}$ is bihomogenous of bidegree $(1,p^{r-1})$, thereby determining a global section of $\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(1,p^{r-1})$.
@@
$JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \frac{m}{p}[p]\ \ \iff\\ JT(\Theta_{\mathfrak{g}/\mathbb{P}^{1},r,(\underline{B},\langle s,t\rangle)},k(\underline{B},\langle s,t\rangle)\otimes M)\ =\ \frac{m}{p}[p].$
@@
As recalled in the proof of Proposition 4.13, $\Theta_{\mathfrak{g},r,\underline{B}}-\Theta_{\mathfrak{g},r,\underline{B}}^{\exp}$ is a linear combination of terms each of which is a product of at least 2 $p$-nilpotent factors. Thus, for any $\langle s,t\rangle\in\mathbb{P}^{1}$, $s\cdot\Theta_{\mathfrak{g},r,\underline{B}}^{\exp}\ +\ t\cdot\Theta_{\mathfrak{g},r,\underline{B}}$ differs from $\Theta_{\mathfrak{g},r,\underline{B}}^{\exp}$ by a sum of terms each of which is a product of at least 2 $p$-nilpotent factors. Consequently, the proposition follows from [8, Prop 8]. ∎
If $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$, if $j$ is a positive integer between 1 and r1r-1, and if $M$ is a finite dimensional $\mathbb{G}_{(r)}$-module, then $M$ is said to have constant $j$-rank if for every scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ the rank of $\Theta_{\mathfrak{g},r,\underline{B}}^{j}:k(\underline{B})\otimes M\to k(\underline{B})\otimes M$ is independent of $\underline{B}$. (See [16, Defn 3.12].) Thus, $M$ has constant Jordan type if and only if $M$ has constant $j$-type for each $j,\ 1\leq j<r$.
@@
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$, and let $M$ be a $\mathbb{G}_{(r)}$-module. Then for any $j,0<j<p$, there are quasi-coherent sheaves ${\mathcal{K}\text{\it er}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\},\ {\mathcal{C}\text{\it oker}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\},\ {\mathcal{I}\text{\it m}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ on $\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ satisfying
@@
    The fibers over $\langle 0,t\rangle\in\mathbb{P}^{1}$ of these sheaves are naturally identified with the sheaves on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ given by the pull-backs along $\Phi_{\mathfrak{g},r}^{*}$ of the corresponding sheaves on $\mathbb{P}V_{r}(\mathbb{G})$ for $\tilde{\Theta}_{\mathbb{G},r,M}^{j}$.
@@
    The fibers over $\langle s,0\rangle\in\mathbb{P}^{1}$ of these sheaves are naturally identified with the corresponding sheaves on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ for $(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}$.
@@
    If $M$ is a finite dimensional $\mathbb{G}_{(r)}$-module of constant $j$-rank, then these sheaves are locally free, coherent sheaves on $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\times\mathbb{P}^{1}$.
In other words, for any $\mathbb{G}_{(r)}$-module of constant $j$-type, taking kernels, cokernels, and images of $(\tilde{\Theta}_{\mathbb{G},r,M})^{j}$ and $(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}$ yield $\mathbb{A}^{1}$-homotopy equivalent vector bundles on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$.
@@
By Proposition 4.5 and Theorem 4.8, $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}$ of Definition 6.5 is bihomogeneous of bidegree $(1,p^{r-1})$. Thus, the bigraded module
(6.7.1) $ker\{(\Theta_{\mathfrak{g}/\mathbb{P}^{1},r})^{j}:(k[s,t]\otimes k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)\ \to\ (k[s,t]\otimes k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)[1,p^{r-1}]$
@@
determines the quasi-coherent sheaves ${\mathcal{K}\text{\it er}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\},\ {\mathcal{C}\text{\it oker}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\},\\ {\mathcal{I}\text{\it m}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ on $\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$. The first two assertions follow from the observation that the fiber above $\langle a,b\rangle\in\mathbb{P}^{1}$ of a sheaf on $\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ is the sheaf associated to the bigraded graded module (6.7.1) obtained by extending along $k[s,t]\to k(\langle a,b\rangle)\in\mathbb{P}^{1}$ which sends $s$ to $a$ and $t$ to $b$.